OpenStudy (anonymous):

Solve: 6(x + 4) = 8(x − 2) A)−4 B)4 C)−20 D)20

OpenStudy (cgreenwade2000):

The first step is to distribute.

OpenStudy (cgreenwade2000):

Dido, your answer is wildly wrong.

OpenStudy (anonymous):

I know. I just realized that.

OpenStudy (cgreenwade2000):

Don't give the answer Dido. Let her see how to solve it.

OpenStudy (anonymous):

Fine.

OpenStudy (anonymous):

no i dont know much about it :(

OpenStudy (anonymous):

Well try to distribute using the distributive property.

OpenStudy (anonymous):

And then simplify both sides.

OpenStudy (anonymous):

Well after distributing you get 6x+24=8x-16 . Can you go from there?

OpenStudy (anonymous):

can you just tell me answer lol

OpenStudy (anonymous):

pweez :3

OpenStudy (cgreenwade2000):

You have to do some of you math homework. Just try it out. It gets easier with practice, but harder if others do it for you.

OpenStudy (anonymous):

i got 30x=-8 so then i divided and came to -3.75 i know thats wrong thats why i need some help with it :c

OpenStudy (cgreenwade2000):

When you distribute, you should get 6x + 24 = 8x - 16 You can add 16 to both sides of the equation ( cancels out the -16 on the right side)

OpenStudy (anonymous):

i subtracted 8 from both sides though

OpenStudy (anonymous):

oh i see

OpenStudy (cgreenwade2000):

Tell me what you get after that step.

OpenStudy (anonymous):

-10+24=8x is that right

OpenStudy (cgreenwade2000):

You cant add 16 + 6x You need to set it up as so: 6x + 24 +16 = 8x -16+16 combine like terms and you should come up with 6x + 40 = 8x (the -16+16 = 0)(24 +16 = 40) Now you tell me the next step.

OpenStudy (anonymous):

what do i do about the 0 do i add the 0 or just leave the 0 out and put 8x